OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

Parth (parthkohli):

\[\cos(3x) = \cos(1.5x + 1.5x)\]Use\[\cos(a + b) = \cos(a)\cos(b) - \sin(a)\sin(b)\]

OpenStudy (anonymous):

2 sin2x cos x - 2 sin x cos x be my answer

OpenStudy (cwrw238):

but that has sin2x in it ??

Parth (parthkohli):

Rather, use \(\cos(2x) = \cos^2 (x) - \sin^2(x)\)

Parth (parthkohli):

Here, \(x = 1.5x\).

OpenStudy (anonymous):

-sin3x + 2 sin x cos x

Parth (parthkohli):

Hmm! I think that doesn't work.

OpenStudy (anonymous):

cos x - 4 cos x sin2x

Parth (parthkohli):

\[\cos(x + 2x) = \cos(x)\cos(2x) - \sin(x)\sin(2x)\]So you can use the formulas for \(\sin(2x)\) and \(\cos(2x)\)

OpenStudy (anonymous):

-sin2x + 2 sin x cos x

OpenStudy (anonymous):

is my answer

Parth (parthkohli):

But that has \(2x\) in it

OpenStudy (anonymous):

These are my choices, one of them has to fit cos x - 4 cos x sin2x -sin3x + 2 sin x cos x -sin2x + 2 sin x cos x 2 sin2x cos x - 2 sin x cos x